Show that the following relations are Partial order relations.

(a) R on the set of integers Z, defined by, R = {(a, b) | a ≤ b}

(b) R on the set of integers Z, defined by, R = {(a, b) | a ≥ b}

(c) R on the set of positive integers N, defined by, R = {(a, b) | a divides b} [Note: It is divisibility

relation]

(d) R on the Power set of a set S, defined by, R = {(A, B) | A is a subset of B} [Note: It is set inclu￾sion relation]


1
Expert's answer
2021-06-29T07:57:25-0400

Let us show that the following relations are partial order relations.


(a) "R" on the set of integers "\Z" , defined by, "R = \{(a, b) | a ≤ b\}."


Since "a\le a" for any "a\in\Z," we conclude that "(a, a)\in R" for any "a\in\Z," and hence the relation "R" is reflexive. If "(a,b)\in R" and "(b,a)\in R," then "a\le b" and "b\le a", and hence "a=b." It follows that "R" is an antisymmetric relation. If "(a,b)\in R" and "(b,c)\in R," then "a\le b" and "b\le c", and thus "a\le c." Therefore, "R" is a transitive relation. We conclude that "R" is a partial order relation.


(b) "R" on the set of integers "\Z" , defined by, "R = \{(a, b) | a ≥ b\}."


Since "a\ge a" for any "a\in\Z," we conclude that "(a, a)\in R" for any "a\in\Z," and hence the relation "R" is reflexive. If "(a,b)\in R" and "(b,a)\in R," then "a\ge b" and "b\ge a", and hence "a=b." It follows that "R" is an antisymmetric relation. If "(a,b)\in R" and "(b,c)\in R," then "a\ge b" and "b\ge c", and thus "a\ge c." Therefore, "R" is a transitive relation. We conclude that "R" is a partial order relation.


(c) "R" on the set of positive integers "\N", defined by, "R = \{(a, b) | a \text{ divides }b\}."


Since "a| a" for any "a\in\N," we conclude that "(a, a)\in R" for any "a\in\N," and hence the relation "R" is reflexive. If "(a,b)\in R" and "(b,a)\in R," then "a| b" and "b| a". Therefore, "a\le b" and "b\le a", and hence "a=b." It follows that "R" is an antisymmetric relation. If "(a,b)\in R" and "(b,c)\in R," then "a| b" and "b| c", and thus "a| c." Therefore, "R" is a transitive relation. We conclude that "R" is a partial order relation.


(d) "R" on the Power set of a set "S", defined by, "R = \{(A, B) | A \text{ is a subset of } B\}."


Since "A\subset A" for any "A\subset S" we conclude that "(A, A)\in R" for any "A\subset S," and hence the relation "R" is reflexive. If "(A,B)\in R" and "(B,A)\in R," then "A\subset B" and "B\subset A", and hence "A=B." It follows that "R" is an antisymmetric relation. If "(A,B)\in R" and "(B,C)\in R," then "A\subset B" and "B\subset C", and thus "A\subset C." Therefore, "R" is a transitive relation. We conclude that "R" is a partial order relation.